OpenStudy (anonymous):

if v1=(2 -5) and v2=(4 -3), then the angle between the 2 vectors is______degrees. Round your answer to two decimal points

OpenStudy (amistre64):

assuming you missed the commas; just work the definition of a dotproduct to solve for alpha:\[|\vec u||\vec v|~\cos\alpha=\vec u \cdot \vec v\]

OpenStudy (amistre64):

not sure if definition is proper usage .. maybe identity?

OpenStudy (anonymous):

I am still confused..

OpenStudy (amistre64):

youll have to be more descriptive of your confusion ... my mind reading skills are subpar.

OpenStudy (anonymous):

lol!

OpenStudy (anonymous):

How do I find the angle based on the two points?

OpenStudy (amistre64):

by working the setup that the dot product gives us.

OpenStudy (amistre64):

do the dot; divide by the lengths, and inverse the cosine

OpenStudy (anonymous):

have you learned dot product (aka inner product)?

OpenStudy (anonymous):

I think, dont you just multiply both of the xs, and ys, and add them together

OpenStudy (anonymous):

yes

OpenStudy (amistre64):

yes, thats the dot process

OpenStudy (amistre64):

lengths and inverse cosine are all thats left to determine

OpenStudy (amistre64):

think of length is the sqrt of a vector dotted to itself ...

OpenStudy (anonymous):

23 is the dot product!

OpenStudy (amistre64):

(2 -5) (4 -3) ------ 8+15 = 23 correct

OpenStudy (anonymous):

Now what do I do, you said divide by length????

OpenStudy (amistre64):

the lengths are therefore just (2 -5) (2 -5) ------ 4+25, sqrt(29) (4 -3) (4 -3) ----- 16+9, sqrt(25)

OpenStudy (amistre64):

yes, lets work some algebra :) \[|\vec u||\vec v|~\cos\alpha=\vec u \cdot \vec v\] \[\cos\alpha=\frac{\vec u \cdot \vec v}{|\vec u||\vec v|}\] \[\alpha=\cos^{-1}\left(\frac{\vec u \cdot \vec v}{|\vec u||\vec v|}\right)\]

OpenStudy (amistre64):

\[\alpha=\cos^{-1}\left(\frac{23}{5\sqrt{29}}\right)\]

OpenStudy (anonymous):

so the angle is the answer of that!!

OpenStudy (amistre64):

keep in mind that the inverse is a radian measure and to convert to degrees may need a 180/pi factor

OpenStudy (amistre64):

yes

OpenStudy (anonymous):

so i do cos^-1(24.77)

OpenStudy (amistre64):

depends on how accurate you need to be, but yes

OpenStudy (anonymous):

I got domain error on my calculator

OpenStudy (amistre64):

your division was wrong :) 23/(5sqrt(29)) is 0.85419...

OpenStudy (anonymous):

oh thank you so much, :))))))) You were so helpful!! the answer is 31.33!!!!

OpenStudy (amistre64):

the domain of the inverse cosine function is -1 to 1 .... hence 24 is not in the domain

OpenStudy (amistre64):

31.33 degrees is correct

OpenStudy (amistre64):

good luck :)

OpenStudy (anonymous):

Thank You!